OpenStudy (anonymous):

Solve the equation for y. 7x – 5y = 10 A. y= 7/5x - 1/2 B.y= 7/5x - 2 C.y= 5/7x - 10/7 D.y= -5x - 7

OpenStudy (haseeb96):

5y=7x-10 y=7/5 x -10/5 y=7/5 x -2 is the correct answer

OpenStudy (anonymous):

Thx

OpenStudy (haseeb96):

please click on best response

OpenStudy (anonymous):

ok